OpenStudy (anonymous):

Enter the solution with the highest value. (2x+3)(x-1)=0

OpenStudy (anonymous):

This must be "Solving Linear Equations" if i'm correct. So, (2x + 3)(x - 1) = 0 2x + 3 = 0 or x - 1 = 0 2x = -3 or x = 1 x = -3/2 or x = 1 highest value of x is 1 x = -(3/2) || x = 1